Solve each of the following systems by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.

2x + 3y = 1
5x + 3y = 16

Answer by ryanman(19)

-(2x + 3y = 1)
5x + 3y = 16
---------------
so
3x=15
x=5
now substitute
5(5) +3y=16
25+3y=16
3y=16-25
3y= -9
y=-3